### v2

# mystyle backgroundcolor=mygray, leftmargin=0.45em, rightmargin=0.45em, innertopmargin=10pt, innerbottommargin=10pt, skipabove=10pt, skipbelow=10pt

Abstract
Software vulnerabilities pose significant risks to the security and integrity of software systems. Prior studies have proposed various approaches to vulnerability detection using deep learning or pre-trained models. However, there is still a lack of detailed explanations for understanding vulnerabilities beyond merely detecting their occurrence, which fails to truly help software developers understand and remediate the issues. Recently, large language models (LLMs) have demonstrated remarkable capabilities in comprehending complex contexts and generating content, presenting new opportunities for both detecting and explaining software vulnerabilities. In this paper, we conduct a comprehensive study to investigate the capabilities of LLMs in both detecting and explaining vulnerabilities, and we propose LLMVulExp, a framework that utilizes LLMs for these tasks. Under specialized fine-tuning for vulnerability explanation, our LLMVulExp not only detects the types of vulnerabilities in the code but also analyzes the code context to generate the cause, location, and repair suggestions for these vulnerabilities. These detailed explanations are crucial for helping developers quickly analyze and locate vulnerability issues, providing essential guidance and reference for effective remediation. We find that LLMVulExp can effectively enable the LLMs to perform vulnerability detection (e.g., achieving over a 90% F1 score on the SeVC dataset) and provide detailed explanations. We also explore the potential of using advanced strategies such as Chain-of-Thought (CoT) to guide the LLMs in concentrating on vulnerability-prone code, achieving promising results.
11footnotetext:  Co-first authors, equally contributed.22footnotetext:  Corresponding author.
I Introduction
A software vulnerability is a flaw or weakness in a system that can be exploited by an attacker to perform unauthorized actions [1, 2, 3, 4, 5]. These vulnerabilities can lead to severe consequences, including data breaches, financial losses, and damage to an organization’s reputation. The increasing complexity and interconnectedness of software systems introduce the great challenge of identifying and mitigating these vulnerabilities effectively.
Current vulnerability detection techniques mainly include pattern based and deep learning based approaches. Pattern based approaches [6, 7] generally rely on manually defined rules to detect vulnerabilities. Deep learning based approaches [1, 2, 8, 9, 10] train the models using existing vulnerability data and various code representation techniques. Despite these advancements, existing methods often fall short of providing detailed explanations of detected vulnerabilities. This lack of robust explanatory capabilities impedes a comprehensive understanding and effective mitigation of vulnerabilities when applied to real-world usage. It is important to propose new techniques that can detect software vulnerabilities and provide additional explanations. Figure 1 shows an example of the vulnerability detection result with and without an explanation which provides comprehensive information for understanding and fixing the vulnerabilities.
In the context of enhancing detection methods, the advent of Large Language Models (LLMs) offers a promising avenue. LLMs, with their advanced generative capabilities, have demonstrated potential in a wide range of applications, such as natural language processing and machine translation. These models can generate extensive textual content and provide contextually relevant information, making them natural candidates for tasks requiring detailed explanations. However, a substantial gap exists between the capabilities of LLMs and the specific requirements of vulnerability detection and explanation. They struggle to effectively detect vulnerabilities and provide accurate explanations without the domain knowledge of vulnerabilities, which requires a deep understanding of code structures, security contexts, and the intricate interplay between various software components. To bridge this gap, it is imperative to fine-tune LLMs specifically for vulnerability detection and explanation. This specialized training aims to endow LLMs with the necessary skills to accurately identify and explain software vulnerabilities.
Figure 1: Example of vulnerability detection result with and without explanation.
The challenges of this fine-tuning process are twofold: obtaining high-quality training data [1, 11, 12, 13, 14] and effectively adapting the models [15, 16, 17]. Curating relevant datasets that encompass diverse and representative software vulnerabilities is crucial. These datasets must capture various vulnerability types, from simple coding errors to complex logic flaws. Furthermore, fine-tuning requires sophisticated techniques to tailor LLMs to the nuanced requirements of vulnerability detection and explanation. Effective fine-tuning can enhance the model’s understanding of code structures and security contexts, thereby improving its detection and explanatory capabilities.
In this study, we explore the enhancement of LLMs for detecting and explaining software vulnerabilities by proposing an automated framework LLMVulExp, which involves leveraging prompt-based techniques to annotate explanatory information for open-source vulnerability data using advanced LLM models. The annotated explanatory information includes the location of the vulnerability and a detailed explanation. Following this, we employ Low-Rank Adaptation (LoRA) [16] fine-tuning methods to refine the LLMs, improving their detection capabilities and establishing a framework for evaluating their explanatory performance [18]. Additionally, we investigate the correlation between the capability of vulnerability detection and explanation, as well as utilize LLMs to follow a Chain-of-Thought (CoT) [19] strategy for concentrated vulnerability detection in key code snippets prone to software vulnerabilities. Furthermore, we address the evaluation of generated vulnerability explanations [20] by proposing new evaluation metrics and an automated review method using LLMs.
Through a comprehensive evaluation of two widely used vulnerability datasets (i.e., SeVC [21] and DiverseVul [11]), we find that LLMVulExp can enable the LLMs to detect vulnerabilities with a high accuracy (e.g., with an F1 score over 90% on SeVC) and provide an effective explanation. We also find that the key code extraction following a CoT strategy can improve the performance of vulnerability detection by a large margin.
Overall, the contributions of this paper are threefold:
  • We present a comprehensive and effective workflow for training, inferring, and evaluating LLMs based on open-source vulnerability data.
  • We pioneer the exploration of the effectiveness of vulnerability explanation models and propose novel evaluation methods and dimensions for these explanations.
  • We analyze the interrelationship between explanatory and detection capabilities, offering valuable insights for the development of explainable vulnerability detection methodologies.
Paper Organization. Section II summarizes the related work. Section III presents the methodology of our study. Section IV discusses the results to our research questions. Section V discusses the implications of our study. Section VI discusses the threats to validity. Section VII concludes the paper.
II Related Work
In this section, we summarize the related work on vulnerability detection and explanation, respectively.
II-A Vulnerability Detection
Prior studies proposed a series of deep learning approaches [1, 2, 8, 9, 10] to detect vulnerabilities. These studies utilized labeled data with and without vulnerability to train neural networks and capture their semantic characteristics.
Apart from deep learning approaches, recent advancements in LLMs have significantly influenced the field of software vulnerability detection. Advanced methodologies such as zero-shot prompting [12, 22], in-context learning [23, 24], and fine-tuning [16, 15, 11] have been employed in vulnerability detection. Cheshkov et al. [20] evaluated the performance of the ChaptGPT and GPT-3 on vulnerability detection tasks and found they failed to classify vulnerable codes in binary and multi-label settings. Gao et al. [25] proposed a benchmark for using LLMs in vulnerability detection and validated that on simple datasets with few-shot prompting, LLMs can achieve performance that is comparable to or exceeds that of deep learning methods. Nong et al. [26] focused on evaluating the performance improvement of vulnerability detection tasks using chain-of-thought prompting. By introducing chain-of-thought forms based on the semantic structure of code, LLMs can achieve higher detection accuracy. Sun et al. [27] further assessed the true reasoning capabilities of LLMs by decoupling their vulnerability reasoning abilities. They found that supplementing LLMs with high-quality vulnerability-related knowledge and contextual information can enhance their performance. Yusuf et al. [28] discovered through experiments that natural language instructions enhance the performance of vulnerability detection tasks across multiple programming languages. Steenhoek et al. [29] surveyed and evaluated eleven LLMs for their capabilities of vulnerability detection, utilizing various types of prompts, including in-context learning and chain-of-thought, which indicates the accuracy limitation of directly applying LLMs to vulnerability detection without finetuning.
These studies underscore the challenges and opportunities of applying LLMs to vulnerability detection. Unlike prior methods, our research focuses on more practical and challenging explanation tasks, and we enhance the vulnerability understanding and analysis capabilities of LLMs through specialized fine-tuning.
II-B Vulnerability Explanation
Although significant research progress has been made in applying deep learning techniques to vulnerability detection, effectively leveraging these techniques for vulnerability explanation remains a challenging issue. Currently, only a limited number of studies focus on the explanatory capabilities of deep learning based models regarding vulnerabilities.
The VulDeeLocator  [30] enhances a Bi-LSTM detector by incorporating an inner multiplication layer, which aids in forecasting vulnerable statements based on the outputs of this layer. In a similar vein, both IVDetect [2] and LineVul [31] develop vulnerability detection models that utilize subgraphs or attention weights derived from the trained detectors to identify vulnerable statements. VELVET [32] prioritizes vulnerable statements by integrating graph-based neural networks with sequence-based neural networks. LineVD [33] addresses statement-level vulnerability detection as a node classification challenge, employing graph neural networks combined with a transformer-based model on PDG for comprehensive learning. VulTeller [34] emphasizes control flows and taint flows to detect more accurate dependencies for localizing vulnerabilities. VulExplainer  [35] is proposed to locate the fine-grained information pertaining to the vulnerability based on graph neural networks. Coca [36] is a dual-perspective contrastive learning enhancement strategy to improve vulnerability explanation methods based on graph neural networks.
The semantic understanding of neural network offers the potential for vulnerability explanation models tailored to software development. However, accurate detection and effective explanation of vulnerabilities using LLMs remain substantial challenges for general-purpose code models. Our work explores the feasibility of fine-tuning specialized LLMs for vulnerability detection and explanation, addressing a crucial gap in this research domain.
III Methodology
Figure 2: Overview of our study.
To address the current gap in generative vulnerability explanation models and enhance the ability of LLMs to detect and analyze software vulnerabilities, we propose a comprehensive framework for fine-tuning and evaluating specialized models for both vulnerability detection and explanation. Figure 2 presents an overview of our framework, namely, LLMVulExp. Specifically, our framework consists of four core stages: ❶ open-source vulnerability data collection, ❷ automated vulnerability explanation annotation based on prompt engineering, ❸ specialized fine-tuning of vulnerability detection and explanation through instruction-based fine-tuning, and ❹ evaluation of generative vulnerability explanation capabilities.
❶ Fine-tuning and Evaluation Vulnerability Dataset Collection: Enhancing the specialized capabilities of LLMs relies on large quantities of high-quality domain-specific data. In the context of vulnerability detection and explanation, the authenticity of the vulnerability code, the diversity of vulnerability types, and the sufficiency of examples of each type are particularly important. In this paper, we conduct the study on two datasets: (1) SeVC [21], which contains four core types and over 50,000 vulnerable code snippets, and (2) DiverseVul [11], which covers 295 real open-source projects and 150 CWE types.
TABLE I: Statistics of the studied datasets.
Dataset Ori. Vul # Ann. Vul # Vul-Type # Eval. Setting
SeVC 56,395 40,491 4 Single/Multi-Type
DiverseVul 18,945 9,161 10 Multi-Type(CWE)
Semantics-based Vulnerability Candidate (SeVC) dataset includes 126 distinct Common Weakness Enumeration (CWE) types of vulnerabilities with 56,395 vulnerable samples and 364,232 non-vulnerable ones. The SeVC dataset is categorized into four primary groups based on the underlying causes of the vulnerabilities: Library/API Function Call, Array Usage, Pointer Usage and Arithmetic Expression. The SeVC dataset includes a significant number of vulnerabilities for each of its four types, making it suitable for fine-tuning and evaluating models designed to detect and explain specific types of vulnerabilities.
DiverseVul is a C/C++ vulnerable source code dataset, which includes 18,945 vulnerable functions and 330,492 non-vulnerable functions derived from 7,514 commits, encompassing 150 CWEs. DiverseVul is currently the largest real-world C/C++ vulnerability dataset, characterized by longer code snippets, coverage of more diverse projects and CWE types, and lower label noise. Therefore, it is used to evaluate the capabilities of LLMs in handling more challenging real-world scenarios involving a broader range of vulnerability types. We select the top ten most frequent CWE types to construct a fine-tuning dataset, ensuring a sufficient number of samples and a well-defined number of classes for our multi-class detection task.
We deduplicate the vulnerability samples using a hash method. Then, we downsample the non-vulnerability samples in a 1:1 ratio to the vulnerability samples to obtain a balanced dataset, aiming to reduce training overhead and avoid model bias. The dataset details are shown in Table 1. We split the processed dataset into training, validation, and test sets in an 80%: 10%: 10% ratio for both vulnerability and non-vulnerability samples to conduct our experiments.
❷ Automated Vulnerability Explanation Annotation: Current open-source vulnerability datasets predominantly encompass information such as source code, vulnerability labels, CWE types, and commit messages. However, they lack detailed explanations of the vulnerability logic within the source code, presenting a significant challenge for vulnerability detection techniques to provide corresponding explanations for the detection results. Manually annotating real-world vulnerable code explanations requires extensive software development experience and in-depth knowledge of software vulnerabilities, which incurs high labor and time costs.
To address this challenge, we propose an automated vulnerability explanation annotation method based on prompt engineering using LLMs. This method leverages the contextual learning and instruction-following capabilities of LLMs, utilizing prompt engineering to achieve large-scale, high-quality automated vulnerability explanation annotation. The prompts decompose the explanation goal into three sub-goals: vulnerability discrimination, code location, and specific explanation. By combining instruction-based prompt templates with well-annotated examples, we stimulate the model’s contextual learning capabilities, ensuring the effectiveness of vulnerability explanation annotation.
In this paper, we use GPT-3.5 [37], accessed via the API provided by OpenAI [38], to implement the annotation process. To address our research questions and experimental needs, we annotated 40,491 and 9,161 vulnerability explanation data points across two datasets, respectively. This effort fills the current gap in vulnerability explanation data.
❸ Specialized Fine-Tuning of Vulnerability Detection and Explanation: The automated explanation annotation of open-source vulnerability data creates a large-scale dataset for vulnerability detection and explanation, addressing data bottleneck issues in the fine-tuning process. To enhance the vulnerability detection and explanation capabilities of LLMs (especially open-source models with lower computational overhead), we fine-tune the general LLMs to enable them to detect and explain specific types of vulnerabilities in real code. We use instruction-based prompts to guide tasks, helping LLMs correctly understand task goals and generate standardized outputs. To minimize the computational overhead of the fine-tuning process, we adopt the parameter-efficient fine-tuning technique LoRA [16], significantly reducing time and space costs.
❹ Evaluation of Generative Vulnerability Explanation: The current lack of research on model-generated vulnerability explanations highlights the need for effective evaluation methods for LLM-generated explanations. Similar to the challenges in annotation, manual evaluation demands significant human and time resources. Additionally, the effectiveness of vulnerability explanations must be assessed across multiple dimensions to ensure they provide valuable assistance to developers in real-world scenarios. To efficiently evaluate vulnerability explanations, we propose an evaluation method based on three dimensions: accuracy, clarity, and actionability. By leveraging prompt engineering, we develop an automated LLM evaluation method. Expert manual verification is employed to validate both the quality of the specialized vulnerability explanation model’s outputs and the feasibility of the LLM-based automated evaluation scheme.
III-A Vulnerability Interpretation Enhancement Prompting
Despite the robust code understanding and analysis capabilities of LLMs, they still face challenges in complex reasoning tasks that require deep understanding of code, strong reasoning abilities, and specialized knowledge of vulnerabilities. These challenges manifest as insufficient detection accuracy and vague vulnerability analysis. To guide LLMs toward better understanding the goals of vulnerability explanation and to enhance their comprehension and analysis of vulnerable code, we have integrated instruction-based fine-tuning techniques with the contextual learning capabilities of LLMs. By leveraging prior knowledge of open-source code vulnerabilities, we design prompt templates for data annotation, fine-tuning, reasoning, and evaluation, effectively empowering the application of LLMs in vulnerability detection and explanation tasks across all crucial stages of the framework.
Figure 3 illustrates the design of our prompt templates. These templates consist of four main components: task description, specific instructions, generation examples, and sample inputs. (1) Task description: Provides a specific template for the current vulnerability detection and explanation, including information on the types of vulnerabilities being detected and the basic input-output format. This helps the LLM understand task requirements and grasp the background knowledge of the vulnerabilities. (2) Specific instructions: Include requirements for the LLM’s input-output format, such as output steps, the range of vulnerability types to focus on, and the length of the output. These requirements leverage the LLM’s instruction-following ability to ensure uniform and standardized output formats, facilitating subsequent content use and analysis. (3) Generation examples: Provide manually screened samples of vulnerable code snippets and effective explanation data pairs to help the annotation model better understand the task and generation goals. (4) Sample input: Used for effectively inputting the code to be explained, and during the annotation stage, it also includes corresponding labels and other relevant information such as CWE, CVE descriptions, or commit messages as supplementary inputs.
Figure 3: Prompting template of Automated Vulnerability Interpretation Labelling.
Through testing and evaluating various vulnerability detection and explanation tasks, we have verified that the prompt templates effectively achieved the expected goals in all stages, resulting in well-performing fine-tuned specialized vulnerability detection and explanation models.
III-B Key Code Extraction Based on Chain-of-Thought (CoT)
One of the major challenges in explaining code vulnerabilities is extracting key code statements closely related to the target vulnerability type from long code snippets. Traditional static analysis methods extract code pattern information through abstract syntax trees, program dependence graphs, or control flow graphs. However, these methods are often difficult to apply in diverse vulnerability detection scenarios. To further enhance the ability of large models to analyze complex code structures, we propose a chain of thought (CoT) enhancement method based on key code extraction. This method leverages prompt engineering to use LLMs for automated key code extraction of vulnerable code. The extracted key code guides the fine-tuning model in locating suspicious vulnerabilities through a thought chain, thereby enabling targeted detection and explanation of the target vulnerabilities.
The extracted key statements are primarily based on the semantic information of the code itself and the type of vulnerability being focused on. After obtaining the key statements for each vulnerability data instance, these key statements are integrated into the prompts used during the model fine-tuning phase in the form of CoT. This approach guides the model step by step to reference the key statements to complete the vulnerability detection, localization, and explanation.
IV Results
In this section, we discuss the results by proposing and answering the following research questions:
  • RQ1: How effective are LLMs in detecting software vulnerabilities?
  • RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
  • RQ3: How do explanations affect the results of vulnerability detection?
  • RQ4: How does the key code extraction impact detection performance?
IV-A RQ1: How effective are LLMs in detecting software vulnerabilities?
Accurate detection is the foundation for correct vulnerability explanation. In this research question (RQ), we first discuss the detection performance of the fine-tuned specialized vulnerability models across various scenarios.
IV-A1 Experimental Setup
Dataset. As discussed in Section III, we select the SeVC [21] dataset and the DiverseVul dataset [11]. Based on the characteristics of the two datasets, we constructed three different vulnerability detection tasks to evaluate the detection performance of LLMVulExp: binary classification vulnerability detection (SeVC), coarse-grained multi-class vulnerability detection (SeVC), and CWE type-based multi-label vulnerability detection (DiverseVul).
Backbone LLMs. We use the Codellama [39] and Llama3 as our backbone LLMs for fine-tuning. CodeLlama is initialized with the weights of Llama2 [40] and fine-tuned on a specialized code dataset, thus possessing strong code understanding and generation capabilities. The Llama-family LLMs have been widely used by prior studies related to LLMs for software engineering [41, 42, 43]. In light of the natural language instruction comprehension capability of the Instruct version and the training cost, we selected CodeLlama-13B-Instruct [44] as our primary model for the experiments discussed in this chapter. For comparison purposes, we also considered the 7B version and the newly released Llama3-8B-Instruct in the discussion section.
Experimental Setting. To comprehensively evaluate the fine-tuned vulnerability LLMs, we design four sub-research questions:
  • (RQ1.1) How effective is the model in detecting a single specific type of vulnerability?
  • (RQ1.2) What impact does the inclusion of other types of vulnerability data in training have on the detection performance of the target type of vulnerability?
  • (RQ1.3) How effective is a unified model in detecting multiple types of vulnerabilities?
  • (RQ1.4) How does the model perform in real-world scenarios with fewer data and more types of vulnerabilities to detect?
To evaluate the accuracy of vulnerability detection, we use Precision, Recall, and F1-Score as our evaluation metrics. For the binary classification scenario in RQ1.1, we directly use these three metrics. For the multi-class classification in RQ1.2, we use Weighted-F1 and Macro-F1. For the multi-label classification in RQ1.3, we use Micro-F1 and Macro-F1. To test the detection accuracy, we selected CodeT5 [45] and CodeBERT [46] as baselines, which have been commonly used by prior studies in software engineering [47, 48]. In particular, we fine-tune them for classification tasks by adding a linear classification layer.
Implementation Details. We implement our approach using the Transformers [17] and PEFT [49] libraries with the PyTorch platform. All experiments are conducted on two NVIDIA A100-SXM4-80GB GPUs platforms, with the token length limit set to 2048. We use the AdamW optimizer and train the models for 3 epochs. For LoRA configuration, we set the learning rate to 0.0003, weight decay to 0.01, the LoRA rank is set to 16, the LoRA scaling factor to 16 and the dropout to 0.05.
IV-A2 RQ1.1: Detection with specified vulnerability type using a dedicated model
To evaluate the detection capability for a single type of vulnerability, we fine-tune each vulnerability type on SeVC to obtain a specialized model for that particular type. The sample size of each type is shown in Table II. In the training and inference prompts, we specify the type of vulnerability to help narrow the scope of detection. During training, we use explanations annotated by GPT-3.5, which include vulnerability location information and specific details, as the target output for the model to perform detection and explanation tasks. The inputs for the model during both training and inference are code snippets from the given examples. As a generative model, we conduct a binary classification task based on the semantics of the generated text. For vulnerability samples, the model generates corresponding explanatory information, while for non-vulnerability samples, it outputs a predefined fixed pattern (e.g., “There are no security issues”).
TABLE II: Performances Comparison of fine-tuning on SeVC(RQ1.1).
Metric API Arith. Pointer Array Average
#Samples 20,294 5,968 38,040 16,680 80,982
Precision (LLMVulExp) 91.6% 90.3% 93.7% 95.3% 92.7%
Precision (CodeLlama) 61.8% 64.5% 70.2% 66.2% 65.7%
Recall (LLMVulExp) 94.5% 93.6% 91.2% 89.7% 92.3%
Recall (CodeLlama) 26.3% 30.4% 50.2% 36.6% 35.9%
F1 (LLMVulExp) 93.0% 91.9% 92.4% 92.4% 92.4%
F1 (CodeLlama) 36.9% 41.4% 58.6% 47.2% 46.0%
TABLE III: Performances of Single-Type Detection on SeVC(RQ1.1).
Precision Recall F1
Type Ours CodeT5 CodeBert Ours CodeT5 CodeBert Ours CodeT5 CodeBert
API 91.6% 92.2% 93.2% 94.5% 88.1% 87.1% 93.0% 90.1% 90.0%
Arithmetic 90.3% 88.2% 90.9% 93.6% 94.7% 98.0% 91.9% 91.3% 91.9%
Pointer 93.7% 93.5% 95.8% 91.2% 95.0% 93.4% 92.4% 94.3% 94.6%
Array 95.3% 92.9% 95.1% 89.7% 94.1% 92.2% 92.4% 93.5% 93.6%
Average 92.7% 91.7% 93.7% 92.3% 93.0% 92.7% 92.4% 92.3% 92.2%
Experimental Results. We first compared the impact of fine-tuning on the detection accuracy of CodeLlama. As shown in Table II, CodeLlama-13B-Instruct lacks precise vulnerability identification capabilities. Fine-tuning significantly improves the detection accuracy. Therefore, in subsequent tasks, we no longer compare with the original CodeLlama. From Table III we can observe that our generative model still achieves comparable performance to other fine-tuned classification models. On the one hand, this demonstrates that our method can effectively enhance the model’s understanding of this type of vulnerability, capturing the key patterns for determining its type. On the other hand, it also reflects that the detection task for a single type of vulnerability is relatively less challenging for LLMs. If the focus of vulnerability types in practice is relatively concentrated, superior performance can be achieved by fine-tuning with data of the target type.
IV-A3 RQ1.2: Detection with specified vulnerability type using an all-in-one model
To explore the impact of training with data from different types of vulnerabilities on the detection effectiveness for target vulnerabilities, we conduct experiments on SeVC by uniformly fine-tuning a model on four types of vulnerabilities and then performing individual detection for each type. Specifically, during training, we use all data from the four types, employing the same prompt as in (RQ1.1) to indicate the vulnerability type of each example for binary classification detection. During inference, we also provide the vulnerability type information for each example. The difference between RQ1.1 and RQ1.2 is that in RQ1.1, we fine-tune a dedicated model for each type of vulnerability, whereas in RQ1.2, we fine-tune an all-in-one model and use prompts to distinguish the type of vulnerability of interest.
TABLE IV: Performances Comparison between Dedicated Mode and All-in-one Mode(RQ1.2).
Accuracy Precision Recall F1
Type Dedicated All-in-one Dedicated All-in-one Dedicated All-in-one Dedicated All-in-one
API 93.0% 97.1% 91.6% 97.4% 94.5% 96.8% 93.0% 97.1%
Arithmetic 91.7% 96.6% 90.3% 97.6% 93.6% 95.6% 91.9% 96.6%
Pointer 92.7% 97.1% 93.7% 97.6% 91.2% 96.6% 92.4% 97.1%
Array 92.7% 96.6% 95.3% 97.0% 89.7% 96.3% 92.4% 96.6%
Average 92.5% 96.9% 92.7% 97.4% 92.2% 96.3% 92.4% 96.9%
Experimental Results. We present the results of the fine-tuned all-in-one model in Table IV. In terms of the performance comparison between the dedicated mode and the all-in-one mode, when given the types of vulnerabilities to be detected in the code, the all-in-one mode consistently achieves significant improvements across various metrics for all types of vulnerabilities. This indicates that the model can learn more patterns related to code vulnerabilities from data that explains multiple types of vulnerabilities, thereby enhancing its general vulnerability detection capabilities. This also reflects that increasing the data volume and the richness of the training data can further improve the performance of vulnerability-specialized LLMs. Moreover, this improvement is due to the ability to avoid confusion and interference that may arise from different types of vulnerabilities when the target type of the code to be detected is specified.
IV-A4 RQ1.3: Detection with an identification of the vulnerability type
In real software development environments, the vulnerability risks faced are often diverse, making it difficult to preemptively sense the potential types of vulnerabilities in the samples to be detected. This requires the large model to accurately identify and classify multiple types of vulnerabilities and provide specific explanations. To test the accuracy of a unified model in this context, we conduct a multi-class vulnerability detection fine-tuning on SeVC with all samples. Specifically, we add a new ‘[type]’ tag to the model’s response output to classify the vulnerability code, including non-vulnerable and the four types of vulnerabilities in SeVC. In the training and inference prompts, we no longer specify the type of vulnerability. Instead, we adopt a multi-type vulnerability detection scenario and provide specific descriptions of the four types of vulnerabilities in the task description.
TABLE V: Performances of Multi-Type Detection on SeVC(RQ1.3).
Precision Recall F1
Type Ours CodeT5 CodeBert Ours CodeT5 CodeBert Ours CodeT5 CodeBert
Non-vul 95.4% 94.7% 94.6% 98.0% 94.6% 96.7% 96.7% 94.7% 95.7%
Array 61.9% 56.0% 58.4% 55.2% 65.6% 62.8% 58.3% 60.4% 60.5%
Pointer 72.9% 72.2% 74.6% 70.9% 66.8% 70.9% 71.9% 69.4% 70.0%
API 48.8% 45.4% 44.5% 46.0% 40.8% 43.5% 47.4% 43.0% 44.0%
Arithmetic 70.7% 65.3% 68.3% 91.6% 88.6% 87.3% 79.8% 75.2% 76.7%
Weighted 79.9% 78.2% 78.9% 80.5% 78.1% 79.1% 80.1% 78.0% 79.1%
Macro 70.0% 66.7% 68.1% 72.4% 71.3% 71.3% 70.8% 68.5% 69.4%
Experimental Results. We present the metrics for each type for the multi-class vulnerability detection task in Table V. Based on the experimental results, we find that: (1) The non-vulnerability type code has a high precision and recall rate, indicating that the model still possesses strong vulnerability code identification capability under multi-type vulnerabilities and can effectively distinguish non-vulnerable code. (2) Compared to the detection results when the type of interest is provided, the overall performance has declined to varying degrees, indicating that the model finds it more challenging to differentiate between the types of vulnerabilities in a multi-type scenario. (3) Specifically, for each type of vulnerability, the Arithmetic Expression performs relatively better than other types. The other three types exhibit a significant degree of confusion and are difficult to identify. This analysis suggests that while the model can effectively identify non-vulnerable code, it struggles with differentiating between various types of vulnerabilities, especially when the data volume for certain types is insufficient. The results highlight the importance of balanced and representative training data for improving the model’s performance across all vulnerability types.
IV-A5 RQ1.4: Identification of the vulnerability type in datasets with more vulnerability types
In real development environments, security risks stem from a wide variety of project types and complex code structures, often involving a broader range of vulnerability models. To better resemble real development conditions, we conduct a multi-label classification task on the top 10 types of CWE in the real-world vulnerability dataset DiverseVul (where a small number of codes have multiple CWE labels). Specifically, we modify the task description to focus on CWE-type vulnerability detection and explanation tasks, adding a ‘[CWE]’ tag to generate a list of CWEs related to the target code. Additionally, we incorporate CWE descriptions as part of the output explanation to help the model better understand the meaning of each CWE type and utilize these descriptions for vulnerability analysis and identification.
TABLE VI: Performances of Multi-Label CWE Detection on DiverseVul (RQ1.4).
Metric CWE ID Overall (Micro)
787 125 119 20 703 416 476 190 200 399 Ours CodeT5 CodeBERT
Precision 79.1% 70.9% 79.3% 68.1% 46.9% 82.1% 74.4% 80.4% 63.6% 88.2% 74.4% 70.8% 65.3%
Recall 83.3% 72.5% 75.8% 71.0% 53.6% 78.0% 56.9% 84.9% 75.0% 83.3% 75.2% 63.1% 61.9%
F1 Score 81.2% 71.7% 77.5% 69.5% 50.0% 80.0% 64.4% 82.6% 68.9% 85.7% 74.8% 67.0% 63.5%
Experimental Results. The experimental results are presented in Table VI. From the results, we find that after fine-tuning, the specialized vulnerability models achieve good detection accuracy for the 10 types of CWE, demonstrating their ability to grasp the pattern information for each CWE type. Unlike the multi-class task in SeVC, the multi-label CWE task does not require the model to differentiate between categories. Each vulnerability type has a more fine-grained and explicit definition, avoiding potential category association issues in SeVC and resulting in better performance. This validates the feasibility of using LLMs for vulnerability detection and explanation in real software development environments, which is illustrated in the following explanation example.
1 Network::FilterStatus Context::onNetworkNewConnection() { 2 onCreate(root_context_id_); 3 if (!wasm_->onNewConnection_) { 4 return Network::FilterStatus::Continue; 5 } 6 if (wasm_->onNewConnection_(this, id_).u64_ == 0) { 7 return Network::FilterStatus::Continue; 8 } 9 return Network::FilterStatus::StopIteration; 10 } {mdframed} [style=mystyle] [label] This function is vulnerable.
[cwe] This function is related to [’CWE-476’]. CWE-476 NULL Pointer Dereference: A NULL pointer dereference occurs when the application dereferences a pointer that it expects to be valid, but is NULL, typically causing a crash or exit.
[location] if (!wasm_->onNewConnection_)
[explanation] The function is vulnerable to a NULL pointer dereference. It checks if ’wasm_->onNewConnection_’ is NULL but does not handle the case when it is NULL. If ’wasm_->onNewConnection_’ is NULL, the function continues its execution, which could lead to a NULL pointer dereference and cause a crash or exit. The vulnerability is related to CWE-476: NULL Pointer Dereference.
TABLE VII: Example of Generated Explanation on DiverseVul.
Summary of RQ1: We find that LLMs are generally effective in detecting vulnerabilities. The effectiveness can be further enhanced when the type of vulnerability in interest is indicated in the prompt.
IV-B RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
In this RQ, we explore the capability of the LLMVulExp in explaining the detected vulnerabilities. Although many evaluation metrics exist in the field of text generation by LLMs, there are no comprehensive criteria for the evaluation of vulnerability explanations. The explanatory information generated by the model is intended to help software developers identify and mitigate potential vulnerability risks, ensuring the accuracy of vulnerability analysis content. Additionally, it should be readable and concise, enabling developers to efficiently acquire the information and provide actionable suggestions for vulnerability remediation.
Considering the characteristics and requirements of the vulnerability explanation task, we propose three evaluation metrics: Accuracy, Clarity, and Actionability, as the evaluation criteria for vulnerability explanations. We use both manual review and automated evaluation by LLMs to assess the explanatory capabilities based on these criteria.
IV-B1 Experimental Setup
In our study, we examine the explanations generated by the fine-tuned model for vulnerability code on the SeVC and DiverseVul test sets, as well as the ground truth annotations based on GPT-3.5. Due to the large scale of the dataset we use and the high evaluation cost, we randomly sample the data based on a 95% confidence level and a 5% confidence interval [50] following prior studies [51, 52, 53] for the fine-tuned model’s correctly classified instances. We obtain 384 and 224 samples from SeVC and DiverseVul, respectively. This sampling method ensures the statistical validity and representativeness of our evaluation.
Our evaluation includes two parts: (1) Manual review: Two authors of this paper, both of whom have extensive research experience in software vulnerabilities and software development, independently conduct manual checks on the samples, scoring them according to the criteria. In cases of disagreement, discussions are held until a consensus is reached. The value of Cohen’s Kappa [54] in this process is 0.76, indicating substantial agreement. This rigorous manual review process ensures the reliability and accuracy of our evaluation. (2) Automated review using LLM: For the LLM automated review, we utilize GPT-3.5, which is instructed in the prompt with a detailed description of the vulnerability explanation evaluation task and each criterion. We ask the LLM to output the scores for three criteria, with 1 indicating satisfaction and 0 indicating non-satisfaction.
This dual approach of manual and automated evaluation allows us to assess the quality of the explanations provided by the fine-tuned model and explore the potential of using LLMs for automated evaluation tasks, which can be valuable for scaling the evaluation process in the future.
Evaluation Metrics. The specific meanings and requirements of the three metrics are as follows:
  • Accuracy: The explanation should correctly identify and describe the vulnerability, ensuring that the provided details are factually correct and relevant to the detected vulnerability.
  • Clarity: The information should be presented in a clear, understandable manner, and structured in a way that facilitates easy comprehension by software developers.
  • Actionability: The explanation should provide actionable suggestions for code modification and remediation, offering guidance on how to fix the identified issue.
IV-B2 Experimental Results
We present the results of our dual evaluation process in Table VIII, the suffix Gen. refers to the explanation generated by LLMVulExp and Ann. refers to the explanation annotated by GPT-3.5. All-Pos. refers to the results that meet the requirements of all of the three metrics. Combining the results from manual and LLM evaluations, we have made the following observations:
(1) Accuracy of Explanations: We find that LLMVulExp generally achieves a high Accuracy (e.g., over 90.0% for SeVC-Gen. for both the manual review and automated review) in the explanation. They essentially pinpoint the code’s vulnerability risks and specific code locations, reflecting that a key factor in the model’s explanatory power lies in the accuracy of vulnerability type identification.
(2) Clarity of Explanations: LLMVulExp achieves a clarity of 81.4% for SeVC and 94.1% for DiverseVul, respectively. The fine-tuned vulnerability explanation model demonstrates feasibility in reducing the barrier for developers to understand and analyze vulnerabilities in practical applications.
(3) Actionability of Explanations: The explanatory generated by LLMVulExp generally includes actionable modification suggestions (i.e., 93.4% for SeVC and 80.5% for DiverseVul), indicating that training for vulnerability explanation can effectively help models provide practical remediation advice for developers.
(4) Effectiveness of Annotation Method: Based on the results of SeVC-Ann. and DIV-Ann., we find that our proposed annotation method can effectively generate explanatory information for vulnerability data. It can stimulate the analytical capabilities of general-purpose large models for software vulnerabilities, reducing the subsequent annotation costs for vulnerability data.
(5) Potential of LLM in Automated Assessment: The automated assessment by the LLM is close to the capabilities of manual review. For SeVC and DiverseVul, the results of All-Pos. comparing manual review and automated review are 76.0% vs. 74.0% and 80.1% vs. 71.9%, respectively. This indicates that LLM has considerable potential in the evaluation of tasks related to the generation of vulnerability-related text. In the future, we can improve the accuracy of evaluations by using a dual verification process of manual and LLM automated reviews, while also reducing manual labor costs.
These observations suggest that the fine-tuned model and the LLM automated review process are valuable tools in the field of vulnerability detection and explanation. The model’s accuracy in identifying vulnerability types is crucial for generating high-quality explanations. The clarity of the explanations is beneficial for practical applications, making it easier for developers to understand and repair vulnerabilities. The effective annotation method can greatly reduce the cost of labeling vulnerability data, which is a significant advantage for scaling up vulnerability analysis. Furthermore, the LLM’s potential in automated assessment could lead to more efficient evaluation processes, balancing the need for accuracy with cost-effectiveness.
TABLE VIII: Distribution (%) of Vulnerability Explanation Review Results (RQ2).
Metric SeVC-Gen. SeVC-Ann. DIV-Gen. DIV-Ann.
Manual
Accuracy 91.1 93.1 73.3 94.1
Clarity 81.4 81.7 94.1 98.6
Construct. 93.4 94.6 80.5 83.2
All-Pos. 76.0 76.0 59.7 80.1
LLM-Automation
Accuracy 90.4 97.6 96.4 96.8
Clarity 74.3 77.2 95.0 96.8
Construct. 83.5 88.6 67.9 71.9
All-Pos. 72.8 74.0 67.9 71.9
1int i; 2for(i = 0; i < 1; i++) 3twoIntsStruct * dataBuffer = (twoIntsStruct *) ALLOCA(100 * sizeof(twoIntsStruct)); 4size_t i; 5for (i = 0; i < 100; i++) 6 dataBuffer[i].intOne = 1; 7 dataBuffer[i].intTwo = 1; 8data = dataBuffer; 9printStructLine(&data[0]); 10 11void printStructLine(const twoIntsStruct * structTwoIntsStruct) { 12 printf($d\n$", structTwoIntsStruct->intOne, structTwoIntsStruct->intTwo); 13} 14delete [] data;
{mdframed}[style=mystyle] [type] pointer
[location] The line ”delete [] data” has a pointer-related issue.
[explanation]
(Analysis:) The issue in this code is related to the ”data” pointer, which is being improperly freed using the ”delete” operator. The ”data” pointer is assigned the address of a memory block allocated on the stack using the ”ALLOCA” function, which allocates memory on the stack instead of the heap. The ”delete” operator is used to free memory allocated on the heap using the ”new” operator, and using it to free memory allocated on the stack may cause undefined behavior or crashes.
(Suggestion:) To fix this issue, the program should use the ”free” function to free memory allocated on the stack, or use the ”new” operator to allocate memory on the heap instead of the ”ALLOCA” function. Additionally, the program should ensure that the ”data” pointer points to a valid memory location before it is used or accessed, and should add proper error handling and validation to ensure that the ”data” pointer behaves correctly and safely in all cases.
TABLE IX: Example of Generated Explanation on SeVC.
IV-B3 Case Study
To more intuitively illustrate the importance of evaluating the quality of vulnerability explanations, we analyze an example generated by our model. As shown in Figure IV-B2, this vulnerability code is sourced from SeVC and pertains to pointer-type vulnerabilities, demonstrating incorrect memory deallocation in C language using the delete operator. The explanation adheres to our output paradigm, generating vulnerability detection, location, and detailed information.
The provided explanation correctly identifies the core issue: the data pointer is allocated on the stack using the ALLOCA function but is improperly freed using the delete operator. The explanation points out that the delete operator should only be used to free memory allocated on the heap via the new operator. Using it to free stack-allocated memory can lead to undefined behavior and potential crashes, meeting the accuracy requirement. Additionally, the explanation offers recommendations for addressing this vulnerability, such as using the free function to release stack-allocated memory or using the new operator for heap allocation, and ensuring pointer safety. This satisfies the actionability requirement. The explanation is clear, comprehensive, logically coherent, and easy to read, meeting the clarity requirement.
During the review, we also observe some instances with insufficient quality of explanations. Issues include failure to identify the main vulnerability, vague location, verbose expression, and lack of remediation suggestions. Hallucinations in explanation generation are relatively rare but primarily involved memorized descriptions of similar code samples from the training set and location confusion. Overall, the fine-tuned model demonstrates a certain capability for generating vulnerability explanations.
Summary of RQ2: We find that LLMVulExp can generate explanations of the vulnerability with high Accuracy, Clarity, and Actionability. We also explore the potential of automated data annotation using LLMs to mitigate the manual effort.
IV-C RQ3: How do explanations affect the results of vulnerability detection?
Considering the detection task as a step in the explanation task, we have conducted a comprehensive evaluation of the detection capabilities of the vulnerability explanation model in RQ1. This leads to an important research question: What is the impact of fine-tuning for vulnerability explanation on the detection capabilities of LLMs compared to fine-tuning specialized detection models? In this RQ, we aim to experimentally explore the influence of explanatory information on the performance of vulnerability detection by comparing the performance of fine-tuned LLMs with and without explanatory information.
Experimental Setup. We conduct our research as ablation studies under the settings of multi-type vulnerability detection scenarios in SeVC (RQ1.3) and multi-label CWE detection scenarios in DiverseVul (RQ1.4). We transform the vulnerability explanation fine-tuning task into a single vulnerability detection task by removing the explanatory information from the training data annotations and modifying the task description and instructions in the prompt. By doing so, we only require the LLMs to output the corresponding vulnerability types and then fine-tune the LLMs to obtain a vulnerability detection model. We compare the detection accuracy of the vulnerability detection model (referred to as Detector) with that of the vulnerability explanation model (referred to as Explainer).
Figure 4: Performance Comparison between Explainer and Detector.
Experimental Results. We present the experimental results for comparison between the Explainer and the Detector on both datasets in Figure 4. The prefix of “Mi” and “Ma” refers to Micro and Macro, respectively. Based on the experimental results, incorporating vulnerability explanation information into the fine-tuning process did not lead to a significant decline in detection performance. For example, the weighted precision on SeVC for the Explainer and Detector are 80.1% and 81.1%, respectively, while 74.8% for both on DiverseVul. In fact, it even resulted in performance improvements under certain metrics on DiverseVul (e.g., the Explainer’s Macro-F1 score is 73.2%, which is nearly 2% higher than that of the Detector.).
These observations indicate that the vulnerability explanation task can coexist with the vulnerability detection task without compromising the model’s accuracy in detecting vulnerabilities. The model’s detection capability remains intact with an additional focus on explanation capabilities. Moreover, the automatically annotated vulnerability explanation data contains sufficient domain-specific knowledge, aiding the model in better understanding and identifying various vulnerability patterns. This suggests that enhancing a model’s explanation capabilities can be achieved without sacrificing detection performance. Furthermore, well-annotated explanatory data can improve the model’s overall understanding and performance in both vulnerability detection and explanation tasks.
Summary of RQ3: Integrating vulnerability explanation into the fine-tuning process does not compromise detection capabilities and may even lead to improved performance for certain vulnerability scenarios.
IV-D RQ4: How does the key code extraction impact detection performance?
In this RQ, we aim to investigate if the LLMs can identify the key code that might be prone to vulnerability and further examine the code with a specific focus. To achieve this, we have conducted annotations of key code extraction on both the SeVC and DiverseVul datasets and fine-tuned the vulnerability explanation LLMs using the key code information.
Experimental Setup. Similar to the process of annotation discussed in Section III, we use GPT-3.5 to extract the key code. To prevent label leakage, the code vulnerability type tags were not visible during the annotation process. For training data that does not output the original code statements, we will nullify its key code information. If it is test data, it will not be included in the evaluation. During fine-tuning and inference, we modify the prompts used, focusing on task description and instruction sections, to guide the model in focusing on the extracted key code for detecting the vulnerability.
TABLE X: F1 Scores of Key Code Extraction on SeVC (RQ4).
Type Single-Type(RQ1.1) Multi-Type(RQ1.3)
W/O Key. With Key. W/O Key. With Key.
API 93.0% 98.7% 79.8% 87.8%
Arith. 91.9% 97.9% 71.9% 96.0%
Pointer 92.4% 99.1% 58.3% 73.5%
Array 92.4% 98.2% 47.4% 77.9%
Average 92.4% 98.5% 64.4% 83.8%
Experimental Results. We present the experimental results of utilizing Key Code Extraction on SeVC in Table X and DiverseVul in Figure 5. We find that there is a noticeable improvement for SeVC (i.e., 98.5% vs 92.4% for Single-Type and 83.8% vs 64.4% for Multi-Type). For DiverseVul, there is a back-and-forth trend comparing the results with key code and without key code. The potential reason might be that the longer and more complex code length makes it difficult to effectively extract the key code on DiverseVul. The results indicate that this enhancement scheme can considerably improve the detection accuracy of the model in different vulnerability explanation tasks. This demonstrates the importance of supplementing code semantic information for the fine-tuning of large vulnerability models and the feasibility of semantic information extraction based on large models. On the one hand, by extracting key portions of the code that are more likely to contain vulnerabilities, the model can concentrate on the most pertinent information, leading to more accurate detection results. On the other hand, key code extraction helps eliminate irrelevant parts of the code, reduce noise, and improve the signal-to-noise ratio. This makes it easier for the model to learn and detect patterns associated with vulnerabilities. This method does not rely on manual feature extraction, making it broadly applicable. Our experimental results preliminarily demonstrate the feasibility of key code extraction through LLMs.
Figure 5: Performances of Utilizing Key Code Extraction on DiverseVul Dataset(RQ4), the numbers in x-axis represent the CWE IDs.
Summary of RQ4: By guiding the LLMs to focus on key code, the performance of vulnerability detection can be improved by a large margin (e.g., from 64.4% to 83.8% for Multi-Type detection on SeVC).
V Discussion
In this section, we discuss the implications of our study.
Implication 1: Broader Applications for Vulnerability-Related Tasks. Our study reveals that integrating vulnerability explanation with detection does not compromise the model’s performance. This finding suggests that we should consider a wider range of vulnerability-related tasks beyond mere detection when fine-tuning LLMs. By doing so, we can potentially enhance the model’s overall understanding and its ability to provide more contextually rich and actionable insights. For instance, tasks such as vulnerability impact assessment, prioritization based on severity, or even automated patch generation could be integrated into the training regime. This holistic approach could lead to the development of more sophisticated tools that not only identify vulnerabilities but also assist in managing and mitigating associated risks.
Implication 2: Effectiveness of LLMVulExp Using Different LLMs. In this paper, we use CodeLlama-13B-Instruct as the primary model to conduct the experiments. To investigate the effectiveness of our framework using different LLMs, we utilize two additional LLMs, CodeLlama-7B-Instruct and Llama3-8B-Instruct, to conduct the experiments following the setting of RQ1.3. Figure 6 shows the results of multi-type vulnerability detection using LLMs of different sizes. We find that LLMVulExp can achieve effective performance with these additional LLMs. Notably, the performance of CodeLlama-13B-Instruct and CodeLlama-7B-Instruct is similar and outperforms Llama3-8B-Instruct. Overall, LLMVulExp demonstrates its effectiveness using different LLMs of various sizes.
Figure 6: Performance of Multi-Type vulnerability detection using LLMs with different sizes.
Implication 3: Dependency on Annotation Quality in Fine-Tuning Frameworks. The effectiveness of fine-tuning heavily depends on the quality of the annotations used for training. Our study highlights the critical role that high-quality annotations play in the fine-tuning process. It is essential to develop robust annotation frameworks that ensure consistency and accuracy. Furthermore, the annotation process itself could be enhanced through the use of semi-automated tools that provide initial labels, which are then reviewed and refined by human experts. This hybrid approach could balance the need for detailed annotations and the practical constraints of manual labeling.
VI Threats to Validity
Construct Validity. We evaluated the explanatory capability using three proposed metrics: Accuracy, Clarity, and Actionability. While these metrics effectively characterize important aspects of vulnerability explanations, they may not comprehensively cover all features. More comprehensive and refined evaluation metrics for vulnerability explanations require further research. To mitigate potential bias, we employed both manual and automated reviews using LLMs. Two authors independently annotated the explanations and resolved any discrepancies until a consensus was reached. The value of Cohen’s Kappa [54] in this process was 0.76, indicating substantial agreement. This rigorous approach helps ensure the reliability and validity of our evaluation.
Internal Validity. Due to the high computational cost of fine-tuning and inference evaluation of LLMs, we were unable to conduct experiments under different data splits and randomization states. The randomness in our experiments (e.g., data splitting, explanation annotation, and fine-tuning process) may affect the results. However, we conducted multiple trials to validate the stability of our conclusions for each experimental setup. Although the number of hyperparameter trials for LoRa configuration and the LLM generation method was relatively limited, and due to GPU constraints, we did not attempt to fine-tune larger models, such as the 34B version, we ensured that these limitations did not fundamentally undermine our primary objective: to explore the potential of conducting vulnerability detection and explanation tasks using LLMs.
External Validity. We chose the advanced open-source code LLM CodeLlama, based on the Llama2 architecture, as our primary model for this research. We tested different versions, including 7B and 13B, and evaluated the performance of the recently released general large language model Llama3 on vulnerability detection and explanation tasks. However, there are many different architectures of general-purpose open-source LLMs and code models, and our experiments were limited to a few models. Additionally, we conducted experiments using two differently configured open-source project-based vulnerability datasets. Although these datasets include a relatively large amount of vulnerable code, there are still differences compared to code in real development environments. Consequently, our experimental conclusions have an inherent risk of not generalizing to other vulnerability datasets or industrial environments. Nonetheless, our careful selection and thorough evaluation process aimed to ensure that our findings remain relevant and insightful within the scope of our study.
VII Conclusion
In this paper, we propose LLMVulExp, a framework designed to detect and explain vulnerabilities using LLMs. The results underscore the potential of LLMs in advancing vulnerability detection and explanation in software security. Our research provides valuable insights into the fine-tuning of LLMs for vulnerability detection and explanation. It highlights the importance of addressing the data volume bottleneck for training vulnerability LLMs in software development. Furthermore, the quality of annotations is paramount for the success of fine-tuning frameworks. Considering these insights, future work can aim to develop more capable and efficient models that significantly contribute to the field of software security.